# The importance of visual modelling languages in generative software engineering

### v1

@@ 1 Introduction
Software engineering (SE) applies engineering principles to the development, operation, and maintenance of software systems, ensuring they are reliable, efficient, and meet user requirements. Generative Artificial Intelligence (GenAI) models, such as pre-trained models Devlin et al. (2019); Vaswani et al. (2017) and large language models (LLMs), are revolutionising fields like computer vision and natural language processing through their ability to generate novel and contextually-appropriate content. In recent times, the interplay between SE and GenAI has been receiving increasing attention. A vast array of applications of pre-trained models and LLMs are surveyed in Huang et al. (2024); Hou et al. (2024). However, when it comes to the type of data used in existing studies, it appears that all studies surveyed focused on text-based datasets, with the most prevalent type of data utilised in training LLMs for SE tasks bring programming tasks/problems expressed in natural language Hou et al. (2024). Similar conclusions are reached in Huang et al. (2024), which focused on 7 sub-tasks of SE: requirements generation, code generation, code summarisation, test generation, patch generation, code optimisation, and code translation. In all cases, the pipeline typically begins with instructions in natural language that need to be used in the context of one or more of these 7 sub-tasks.
@@
The rest of this paper is organised as follows: Section 2 provides relevant background in GenAI and SE. Section 3 surveys contemporary academic literature on role of GenAI in SE. Section 4 investigates the role of multimodal GPTs in software development by illustrating a selection of use cases. Section 5 provides some concluding remarks by reflecting on how the former use cases address several research gaps in Generative Software Engineering (GenSE), as well as some long standing issues in SE.
@@ 2 Background
2.1 Software Engineering
The term “software engineering” was coined by Margaret Hamilton, who developed on-board flight software for NASA’s Apollo program. Key milestones in the field include the development of the Waterfall model (1970s), the introduction of Object-Oriented programming (1980s), the Agile Manifesto (2001), which advocates for flexible, iterative development, the rise of DevOps (2000s) and of continuous integration/continuous delivery (CI/CD) pipelines (2010s).
In the original Waterfall model, the software development process proceeds through a series of sequential phases: requirement analysis; software design; software development; software quality assurance; and software maintenance. To address the inflexibility of the Waterfall model, the increasing complexity of software, and poor customer satisfaction, in the early 2000s a set of frameworks and practices were proposed based on the values and principles expressed in the Manifesto for Agile Software Development Beck et al. (2001). Core values and principles of Agile include: favouring individuals and interactions over processes and tools; working software over comprehensive documentation; customer collaboration over contract negotiation; and responding to change over following a plan. Agile methodologies promote iterative development, frequent feedback, and adaptability to changing requirements throughout the aforementioned phases.
At the same time, the Unified Modelling Language (UML) emerged in 1990s as a response to the disparate notational systems and approaches to software design Rumbaugh et al. (2004) utilised across the phases of the software development process. UML is a general-purpose standardised visual language for designing systems. In SE, UML is utilised in three main areas: use case development, static analysis, and dynamic analysis of the software system. Use case development (or development of user stories in Agile) is a key step of requirement analysis. Use cases describe how a user interacts with a system or product to achieve a specific goal. Static analysis of a software system describes the structure of a system by showing its classes, their attributes, operations (or methods), and the relationships among objects. Dynamic analysis expresses and model the behaviour of the system over time. More specifically, in UML, use case diagrams support use case development. Class diagrams support static analysis; and interaction (sequence, activity, collaboration) diagrams support dynamic analysis. However, it should be noted that UML is a vast language that finds countless applications to go beyond the three areas here considered.
Despite UML being the standard visual language for designing systems, recent studies suggest that that most software developers favour informal hand drawn diagrams and do not use UML; those using UML, tend to use it informally and selectively Baltes and Diehl (2014). The authors further advocated the development of suitable tools to make better use of such sketches. Our study argues that GenAI may represent such missing tool.
@@ 2.2 Generative Artificial Intelligence
GenAI refers to a class of AI models that can create new content, such as text, diagrams, or code, based on the patterns they learn from existing data Eapen et al. (2023). Large Language Models (LLMs) are a specific category of GenAI models that focus on language-related tasks, such as text generation, translation, summarisation, and question answering. LLMs are typically based on a neural network architecture called a Transformer Vaswani et al. (2017), which is Pre-trained — hence the name Generative Pretrained Transformer (GPT) — on a massive dataset of text and code, and which allows them to process and generate text by considering long-range dependencies and contextual information Brown et al. (2020). The latest GPTs, such as GPT-4 OpenAI (2024), are multimodal: they accept image and text inputs and produce text outputs, leading to new applications Li et al. (2024b).
2.2.1 Prompt engineering
@@
Prompts are functions that map an example from a dataset to a natural language input and target output Bach et al. (2022). Due to the cost of re-training LLMs, prompt engineering, that is the use of prompts to train and query language models, offers a cost-effective way to adapt pre-trained models without full fine-tuning. Key prompt engineering techniques include: Single Prompt Techniques, Multiple Prompt Techniques, and the use of LLM frameworks with external tools.
Single Prompt Techniques include Zero-Shot, Few-Shot, Chain of Thought (CoT), and Program-Aided Language Models (PAL). Zero-Shot Prompting involves providing tasks with natural language without additional context, relying on the model’s pre-existing capabilities. Few-Shot Prompting includes providing examples within the prompt to guide the model, enhancing its performance on complex tasks by exposing it to the input and output patterns Brown et al. (2020). Chain of Thought Prompting Wei et al. (2024) aids in breaking down reasoning tasks into smaller steps to improve processing and outcome accuracy, using either zero-shot or few-shot methods to encourage step-by-step thinking. PAL extends this by integrating code into reasoning processes to produce more accurate outputs Gao et al. (2023).
Multiple Prompt Techniques leverage combining different strategies to refine outputs. Voting/self-consistency Wang et al. (2023) involves generating multiple responses and selecting the most common result, which can improve accuracy, especially for complex reasoning tasks. Divide and Conquer methods split tasks into subtasks handled in sequence for improved manageability and precision, seen in Directional Stimulus Prompting Li et al. (2024a), Generated Knowledge Liu et al. (2021), and Prompt Chaining. Self-evaluation asks the model to verify output accuracy, exemplified by Reflexion Shinn et al. (2024) and Tree of Thoughts Yao et al. (2024), enabling iterative improvement.
Retrieval-Augmented Generation (RAG) Fan et al. (2024) and ReAct Yao et al. (2023) combine LLMs with external systems to improve context handling and output relevance. RAG uses vector search for context retrieval before passing it to an LLM for output generation. ReAct combines reasoning traces and task-specific actions, allowing models to interact with external knowledge sources to refine outputs.
3 The role of GenAI in Software Engineering
@@
In recent times, SE has become one of the important application areas for GenAI. We focus on two recent surveys Huang et al. (2024); Hou et al. (2024) investigating the interplay between SE and GenAI.
@@
Several studies, e.g. Arora et al. (2024); White et al. (2024), investigated the use of GenAI in the context of the requirement engineering sub-task. We do not believe sketches and diagram play a significant role in this sub-task, which is mainly concerned in turning requirements expressed in natural language by the customer into suitable user stories and/or conceptual diagrams Robeer et al. (2016).
SE sub-tasks of interest in the context of the present study include: software design, software development, and code summarisation.
@@
Application of LLMs in software design remains relatively sparse: Huang et al. (2024) does not cover this sub-task, while Hou et al. (2024) only report 4 works Kolthoff et al. (2023); Mandal et al. (2023); White et al. (2024); Zhang et al. (2024), none of which overlaps with the content of the present study; they also stress that by expanding the use of LLMs to this under-explored area, it is possible to improve how software designs are conceptualised.
@@
Conversely, both surveys identify a plethora of works concerned with software development (including code generation, test case generation, patch generation, and code optimisation) and code summarisation.
@@
Code generation has been object of investigation for a long time in the AI community. Early works used symbolic and neural-semiotic approaches Alur et al. (2013). However, recent neurolinguistic models, such as GPT-4 Liu et al. (2024) and Copilot Ma et al. (2023), can generate code directly from natural language descriptions. This capability has significantly advanced code completion, automatic code generation, and the transformation of natural language annotations into code, empowering developers with powerful tools and driving automation in software development. None of the studies listed in the above surveys focus on code generation that leverages multimodal prompts that include UML diagrams.
@@
Code summarisation Ahmed et al. (2024) aims to automatically generate descriptions of a given source code. This technique improves code comprehension, documentation, and collaboration by providing clear summaries. Existing studies in code summarisation focus on analysing code structures and contexts to generate informative natural language summaries. None of the studies listed in the above surveys focus on code summarisation producing a UML diagram as its output.
@@
Finally, while there exist a few studies that investigated the generation of UML diagram with support from LLMs Conrardy and Cabot (2024); Wang et al. (2024); Cámara et al. (2023), none of these studies go as far as investigating the generation of working code from UML diagrams, as well as the reverse engineering of relevant UML diagrams from existing code. Perhaps the most interesting study among the three listed is Cámara et al. (2023), which focus on how to build UML class diagrams by using ChatGPT as a modelling assistant by leveraging the PlantUML language, which we will also adopt in the following sections.
@@ 4 Multimodal GPTs in software development
In what follows, we will assume that a preliminary requirement analysis has been carried out, which has produced a portfolio of initial user stories. While GPTs may find applications in the realm of the development of user stories and in the translation of user stories into working software, as things stand today, this level of automation in the realm of code generation does not exist in SE practice. Conversely, what we find in practice are teams of software developers tasked with translating user stories into working software. Our focus in this section, is to explore novel use cases of GPTs in this specific context.
The role of sketches and diagrams in the daily work of software developers has been investigated in Baltes and Diehl (2014). This study found that most practitioners produced informal hand drawn diagrams and did not use UML; those using UML, tended to use it informally and selectively. We argue that this stems from the fact that the adoption of a “formal” notation, at present, bears no advantage with respect to an “informal” one. Over several decades, firms have repeatedly tried to develop tools (e.g. IBM Rational Rose) that could automatically generate working code from UML diagrams, or reverse engineer UML diagram from existing code. While these technologies still exist, it is rare to find developers who routinely develop a complete set of UML diagrams and then translate this to code using such tools. We find it plausible that the lacklustre success of these tools is due to a fundamental misunderstanding of what UML is: a language for capturing and communicating conceptual requirements, not one for generating complete code. In other words, UML diagrams are most useful when they are high level and sufficiently abstract. Generation of complete working code requires such diagrams to reach a level of detail that is equivalent to the working code itself; but generating diagrams at this level of abstraction would be a complete waste of time: why then not generating the code itself directly?
Albeit we believe there is a good reason for practitioners to produce informal hand drawn diagrams, we also believe that this practice is set to change with the advent of multimodal GPTs, such as GPT-4 OpenAI (2024). In what follows, we will outline a set of use cases illustrating how practitioners may combine multimodal GPTs and UML diagrams to innovate SE practices. We will utilise Microsoft Copilot in its web-based version, which is based on GPT-4o and allows image attachments.
@@ 4.1 Static modelling
As a motivating example to illustrate our use cases, we consider the introductory case study in Section 3 of Stevens and Pooley (2006). In this section, we focus on static modelling; in particular, we assume that the developer has already converted the relevant requirements into a preliminary class diagram such as that shown in Figure 1.
Figure 1: Library class model
The developer now wants to obtain a preliminary code that matches class diagram. Rather than using a UML diagram to code conversion tool, we will here rely on Copilot prompting with images. To achieve this, we attach Figure 1 and use the prompt illustrated in the following greyed box.
@@
The resulting code is presented in our supplementary material (SM). Of course, depending on the specific prompt and GPT system used, the resulting code may vary. These variations are irrelevant, as SE typically proceeds via successive refinements to achieve the desired outcome.
@@ 4.2 Comparison to natural language
Figure 4: The description of the class diagram in PlantUML notation as returned by Copilot
This example makes apparent the point discussed in the paragraph from Tufte (2003) we quoted in Section 1: the visual channel is an intensely high-resolution channel.
@@ 4.3 Dynamic modelling
Figure 5: A sequence diagram
@@
The resulting code is presented in our SM.
@@
The resulting code is presented in our SM.
@@
Unfortunately, the code obtained is not fully functional. There is a bug in method borrow_copy of class LibraryMember. This specific bug can be easily resolved by adding a single instruction (more details are provided in the SM). In general, further iterations of the process, potentially involving additional diagrams and a test-driven development strategy, can be used to obtain working code. In fact, the specific code generated or the absence/presence of bugs are both immaterial to the case study here presented, which is simply meant to illustrate potential use cases that combine multimodal GPTs and UML diagrams to innovate and support SE practices.
4.4 Hand drawn diagrams

### v2

@@ 1 Introduction
Software engineering (SE) applies engineering principles to the development, operation, and maintenance of software systems, ensuring they are reliable, efficient, and meet user requirements. Generative Artificial Intelligence (GenAI) models, such as pre-trained models Devlin et al. (2019); Vaswani et al. (2017) and large language models (LLMs), are revolutionising fields like computer vision and natural language processing through their ability to generate novel and contextually-appropriate content. In recent times, the interplay between SE and GenAI has been receiving increasing attention. A vast array of applications of pre-trained models and LLMs are surveyed in Huang et al. (2024); Hou et al. (2024). However, when it comes to the type of data used in existing studies, it appears that all studies surveyed focused on text-based datasets, with the most prevalent type of data utilised in training LLMs for SE tasks bring programming tasks/problems expressed in natural language Hou et al. (2024). Similar conclusions are reached in Huang et al. (2024), which focused on seven sub-tasks of SE: requirements generation, code generation, code summarisation, test generation, patch generation, code optimisation, and code translation. In all cases, the pipeline typically begins with instructions in natural language that need to be used in the context of one or more of these seven sub-tasks.
@@
The rest of this paper is organised as follows: Section 2 provides relevant background in GenAI and SE. Section 3 outlines the methodology adopted in our study. Section 4 investigates the role of multimodal GPTs in software development by illustrating a selection of use cases. Section 5 provides concluding remarks by reflecting on how the former use cases address several research gaps in Generative Software Engineering (GenSE), as well as some longstanding issues in SE, and open new research directions.
@@ 2 Background
2.1 Software Engineering and UML
The Unified Modelling Language (UML) emerged in 1990s as a response to the disparate notational systems and approaches to software design Rumbaugh et al. (2004) utilised across the phases of the software development process. UML is a general-purpose standardised visual language for designing systems. In SE, UML is utilised in three main areas: use case development, static analysis, and dynamic analysis of the software system. Use case development is a key step of requirement analysis. Use cases [^1] User stories in Agile Software Development Beck et al. (2001) describe how a user interacts with a system or product to achieve a specific goal. Static analysis of a software system describes the structure of a system by showing its classes, their attributes, operations (or methods), and the relationships among objects. Dynamic analysis expresses and model the behaviour of the system over time. More specifically, in UML, use case diagrams support use case development. Class diagrams support static analysis; and interaction (sequence, activity, collaboration) diagrams support dynamic analysis. However, it should be noted that UML is a vast language that finds countless applications to go beyond the three areas here considered.
Despite UML being the standard visual language for designing systems, recent studies suggest that that most software developers favour informal hand-drawn diagrams and do not use UML; those using UML, tend to use it informally and selectively Baltes and Diehl (2014). The authors further advocated the development of suitable tools to make better use of such sketches. Our study argues that GenAI may represent such missing tool.
@@ 2.2 Generative Artificial Intelligence
GenAI refers to a class of AI models that can create new content, such as text, diagrams, or code, based on the patterns they learn from existing data Eapen et al. (2023). Large Language Models (LLMs) are a specific category of GenAI models that focus on language-related tasks, such as text generation, translation, summarisation, and question answering. LLMs are typically based on a neural network architecture called a Transformer Vaswani et al. (2017), which is pre-trained — hence the name Generative Pretrained Transformer (GPT) — on a massive dataset of text and code, and which allows them to process and generate text by considering long-range dependencies and contextual information Brown et al. (2020). The latest GPTs, such as GPT-4 OpenAI (2024), are multimodal: they accept image and text inputs, and produce image and text outputs, leading to new applications.
@@
Prompts are instructions or input given to a large language model (LLM) to generate a specific response. Prompt engineering, that is the design prompting strategies to query language models, offers a cost-effective way to adapt pre-trained models without full fine-tuning. Prompt engineering techniques include: Single Prompt Techniques (e.g. Zero-Shot, Few-Shot), Multiple Prompt Techniques (e.g. Prompt Chaining), and the use of LLM frameworks with external tools (e.g. Retrieval-Augmented Generation Fan et al. (2024), and ReAct Yao et al. (2023)).
2.3 The role of GenAI in Software Engineering
@@
In recent times, SE has become one of the important application areas for GenAI. We focus on two recent surveys Huang et al. (2024); Hou et al. (2024) investigating the interplay between SE and GenAI over a large body of recent works.
@@
Several studies, e.g. Arora et al. (2024); White et al. (2024), investigated the use of GenAI in the context of the requirement engineering sub-task. GenAI does play a role in this sub-task, which is mainly concerned in turning requirements expressed in natural language by the customer into suitable user stories and/or conceptual diagrams Robeer et al. (2016), but our focus in this study will not be on this sub-task. Conversely, SE sub-tasks of interest in the context of the present study include: software design, software development, and code summarisation.
@@
Application of LLMs in software design remains relatively sparse: Huang et al. (2024) does not include any study within this sub-task; while Hou et al. (2024) only report 4 works Kolthoff et al. (2023); Mandal et al. (2023); White et al. (2024); Zhang et al. (2024), none of which overlaps with the content of the present study; they also stress that by expanding the use of LLMs to this under-explored area it is possible to improve how software designs are conceptualised.
@@
Both surveys identify a plethora of works concerned with software development (including code generation, test case generation, patch generation, and code optimisation) and code summarisation.
@@
Code generation has been object of investigation for a long time in the AI community. Early works used symbolic and neural-semiotic approaches Alur et al. (2013). However, recent neurolinguistic models, such as GPT-4 Liu et al. (2024) and Copilot Ma et al. (2023), can generate code directly from natural language descriptions. This capability has significantly advanced code completion, automatic code generation, and the transformation of natural language annotations into code, empowering developers with powerful tools and driving automation in software development. None of the studies listed in the above surveys focus on code generation leveraging multimodal prompts that include sketches & diagrams.
@@
Code summarisation Ahmed et al. (2024) aims to automatically generate descriptions of a given source code. This technique improves code comprehension, documentation, and collaboration by providing clear summaries. Existing studies in code summarisation focus on analysing code structures and contexts to generate informative natural language summaries. None of the studies listed in the above surveys focus on code summarisation producing a diagram as its output.
@@
Finally, while there exist a few studies that investigated the generation of UML diagram with support from LLMs Conrardy and Cabot (2024); Wang et al. (2024); Cámara et al. (2023), none of these studies go as far as investigating the generation of working code from UML diagrams, as well as the reverse engineering of relevant UML diagrams from existing code. Perhaps the most interesting study among the three listed is Cámara et al. (2023), which focuses on building UML class diagrams in PlantUML notation, which we also adopt in this work, by using ChatGPT as a modelling assistant prompted with instructions in natural language.
3 Methodology
Our approach takes the form of an exploratory case study Priya (2020), which aims to identify new research questions that can be investigated in future studies. More specifically, we develop a portfolio of novel GenSE use cases that, to the best of our knowledge, have not been previously investigated in the literature.
In our study, we utilise Microsoft Copilot in its web-based version, which is based on GPT-4o and allows image attachments. To ensure reproducibility of the discussion in Sections 4.1-4.5, we have also developed an equivalent pipeline in Gemini, formalised in a Jupyter Notebook based on gemini-1.5-flash, which is included in the supplementary material (SM). In both cases, we left all LLM parameters to their default settings and we did not specify any role or context instructions.
Table 1 maps use cases discussed in the rest of this work to relevant SE sub-tasks of interest. All use cases are based on an interaction with the LLMs that takes the form of an interactive chat comprising multiple rounds of questions and responses, which allow users to step incrementally toward answers and thus get help with multipart problems. The core principle underpinning all our use cases is to leverage UML diagrams in order to guide the software development process. More specifically, in Section 4.1 we leverage class diagrams to guide the LLM in the context of implementing relevant classes, attributes, and operations; in Section 4.2, we contrast diagram-based and natural language-based descriptions of a system; in Section 4.3 we leverage interaction diagrams to guide the LLM in the context of implementing the desired system behaviour; in Section 4.4 we leverage hand drawn activity diagrams to guide the LLM in the context of implementing a given method; in Section 4.5 we leverage design patterns to influence the design of a given system. Finally, in our SM we present three additional case studies: in the first case study, we leverage hand drawn diagrams and design patterns to implement a mathematical expressions evaluator; in the second and third case studies, which feature a higher degree of complexity, we leverage class diagrams to implement a tic tac toe game and a game of checkers, respectively.
Use case Section SE sub-tasks
static modelling 4.1 SD/DE/CS
information transfer 4.2 SD
dynamic modelling 4.3 SD/DE/CS
hand drawn diagrams 4.4 DE/CS
design patterns 4.5 SD
expression evaluator SM SD/DE
tic tac toe SM SD/DE
checkers SM SD/DE
Table 1: Use cases investigated in the rest of this work (SD: software design; DE: software development; CS: code summarisation)
@@ 4 Multimodal GPTs in software development
In what follows, we will assume that a preliminary requirement analysis has been carried out, which has produced a portfolio of initial user stories. While GPTs may in principle support automated elicitation of user stories and fully automated translation of user stories into working software, as things stand today this level of automation in the realm of code generation is hardly found in SE practice; nor we suspect it would be beneficial, as it may conflict with some of the principles of Agile. What we find in practice are teams of software developers collaborating to translate user stories into working software. Our focus is to illustrate novel use cases of GPTs in this specific context.
The role of sketches and diagrams in the daily work of software developers has been investigated in Baltes and Diehl (2014). This study found that most practitioners produced informal hand-drawn diagrams and did not use UML; those using UML, tended to use it informally and selectively. We argue that this stems from the fact that the adoption of a “formal” notation, at present, bears no advantage with respect to an “informal” one. Over several decades, firms have repeatedly tried to develop tools (e.g. IBM Rational Rose) that could automatically generate working code from UML diagrams, or reverse engineer UML diagram from existing code. While these technologies still exist, it is rare to find developers who routinely develop a complete set of UML diagrams and then translate this to code using such tools. We find it plausible that the lacklustre success of these tools is due to a fundamental misunderstanding of what UML is: a language for capturing and communicating conceptual requirements, not one aimed at describing a complete system. In other words, UML diagrams are most useful when they are high level and sufficiently abstract. Generation of complete working code requires such diagrams to reach a level of detail that is equivalent to the working code itself; but generating diagrams at this level of abstraction would be a complete waste of time: why then not generating the code itself directly?
Albeit there are already good reasons for practitioners to produce sketches and diagrams in certain circumstances, we believe the advent of multimodal GPTs, such as GPT-4 OpenAI (2024), will substantially increase the associated use cases. To exemplify this, in what follows we outline a set of use cases illustrating how practitioners may combine multimodal GPTs and UML diagrams to innovate SE practices.
@@ 4.1 Static modelling
As a motivating example to illustrate our use cases, we consider the introductory case study in Chapter 3 of Stevens and Pooley (2006). In this section, we focus on static modelling; in particular, we assume that the developer has already converted the relevant requirements into a preliminary class diagram such as that shown in Figure 1.
Figure 1: Class model of a library
The developer now wants to obtain a preliminary code that matches this class diagram. Rather than using a UML diagram to code conversion tool, we will here rely on Copilot prompting with images. To achieve this, we attach Figure 1 and use the prompt illustrated in the following greyed box.
@@
The resulting code is presented in the SM. It is noteworthy that in addition to implementing the classes, Copilot also made an attempt at drafting part of the business logic by leveraging semantic information associated with the labels of the relationships in the diagram.
@@ 4.2 Comparison to natural language
Figure 4: The description of the class diagram in PlantUML notation
This example makes apparent the point discussed in the paragraph from Tufte (2003) we quoted in Section 1: the visual channel is an intensely high-resolution channel. Moreover, it reinforces the findings in Li et al. (2024) on the interrelation between understanding, learning, and compression. Further examples contrasting diagrams and natural language descriptions are provided in the SM.
@@ 4.3 Dynamic modelling
Figure 5: A sequence diagram illustrating an interaction
@@
The resulting code is presented in the SM.
@@
The resulting code is presented in the SM.
@@
Unfortunately, the code obtained is not fully functional. There is a bug in method borrow_copy of class LibraryMember. This bug can easily be resolved by adding a single instruction (more details are provided in the SM). In line with contemporary SE and Agile practices — which predicate an iterative form of software development aimed at progressively discovering how to meet customer requirements — further iterations of the process, potentially involving additional diagrams and further prompts, can be used to obtain working code.
4.4 Hand-drawn diagrams